# A lower bound on the number of edges in DP-critical graphs

### v1

@@ Abstract
1 Introduction
Let $\mathbb{N}$ denote the set of positive integers. For a set $X$, let $\operatorname{Pow}(X)$ denote the power set of a set $X$, and denote $\bigcup_{v\in X}f(v)$ by $f(X)$. A graph is simple if it has no parallel edges. Given a multigraph $G$, let $V(G)$ and $E(G)$ denote the vertex set and edge set, respectively, of $G$. Let $G^{-}$ denote a multigraph obtained from $G$ by removing any edge. Given a vertex $v\in V(G)$, the degree of $v$, denoted $d_{G}(v)$, is the number of edges incident with $v$. The neighborhood $N_{G}(v)$ of a vertex $v$ is the set of vertices adjacent to $v$. Note that $|N_{G}(v)|\leq d_{G}(v)$, and equality holds if and only if $v$ has no incident parallel edges. A vertex of degree $d$ (at least $d$) is a $d$-vertex ($d^{+}$-vertex). For vertex subsets $S_{1}$ and $S_{2}$, let $E_{G}(S_{1},S_{2})$ denote the set of edges $xy\in E(G)$ where $x\in S_{1}$ and $y\in S_{2}$.
1.1 Known results on proper coloring
A (proper) $k$-coloring of a graph $G$ is a mapping $g\,:\,V(G)\to\{1,\ldots,k\}$ such that $g(u)\neq g(v)$ for each $uv\in E(G)$. The chromatic number of $G$, denoted $\chi(G)$, is the minimum positive integer $k$ for which $G$ has a proper $k$-coloring. A graph $G$ is $k$-colorable if $\chi(G)\leq k$. For a positive integer $k$, a graph $G$ is $k$-critical if $\chi(G)=k$, but every proper subgraph of $G$ is $(k-1)$-colorable.
The notion of $k$-critical graphs was introduced and systematically studied by Dirac [6, 7, 8, 9, 10, 11, 12] as early as 1951. In particular, Dirac considered the minimum number $f(n,k)$ of edges in an $n$-vertex $k$-critical graph. With this notation, $f(k,k)=\binom{k}{2}$ and $f(k+1,k)$ is not well defined. Already simplest bounds on $f(n,k)$ yield the Heawood Formula [25, (1.1)] that every graph $G$ embeddable on an orientable surface $S_{\gamma}$ of genus $\gamma\geq 1$ satisfies $\chi(G)\leq\lfloor c_{\gamma}\rfloor$, where $c_{\gamma}=\frac{7+\sqrt{1+48\gamma}}{2}$. Dirac sharpened this result by showing that for $\gamma\geq 1$ every graph embeddable on $S_{\gamma}$ with chromatic number $\lfloor c_{\gamma}\rfloor$ contains the complete graph with $\lfloor c_{\gamma}\rfloor$ vertices as a subgraph. For this result, he used the following lower bound on $f(n,k)$:
@@ 1.1 Known results on proper coloring
$f(n,k)\geq\left(k-1\right)\frac{n}{2}+\frac{k-3}{2}.$ (1)
@@
In his fundamental papers [15] and [16] from 1963, Gallai proved a series of important properties of color-critical graphs. Recall that a Gallai tree is a connected graph whose every block is a complete graph or an odd cycle. A Gallai forest is a graph where every connected component is a Gallai tree.
@@
Let $k\geq 4$, and let $G$ be a $k$-critical graph. If $B$ is the set of $(k-1)$-vertices in $G$, then $G[B]$ is a Gallai forest.
Theorem 1.2 implies the following lower bound on $f(n,k)$:
Theorem 1.3 (Gallai [15]).
If $k\geq 4$ and $n\geq k+2$, then
$f(n,k)\geq\left({k-1}+\frac{k-3}{k^{2}-3}\right)\frac{n}{2}.$ (2)
For large $n$, this bound is much stronger than the bound in Theorem 1.1.
Krivelevich [21, 22] improved the lower bound on $f(n,k)$ in Theorem 1.3 to
$f(n,k)\geq\left(k-1+\frac{k-3}{k^{2}-2k-1}\right)\frac{n}{2}$ (3)
and demonstrated nice applications of his bound. Then Kostochka and Stiebitz [20] proved that for $k\geq 6$ and $n\geq k+2$,
$f(n,k)\geq\left(k-1+\frac{2(k-3)}{k^{2}+6k-9-6/(k-2)}\right)\frac{n}{2}.$ (4)
Note that the last term in the bound (4) is asymptotically (in $k$) twice larger than the one in (2). Then, Kostochka and Yancey [18] proved an asymptotically exact bound:
Theorem 1.4 (Kostochka and Yancey [18]).
If $k\geq 4$ and $n\geq k,\,n\neq k+1$, then
$f(n,k)\geq\left(k-1+\frac{k-3}{k-1}\right)\frac{n}{2}-\frac{k(k-3)}{2(k-1)}.$ (5)
In this bound, the last term in the coefficient of $\frac{n}{2}$ is asymptotically (in $k$) $k$ times larger than the one in (2).
1.2 Known results on list coloring
List coloring was introduced independently by Vizing [26] and Erdős, Rubin, and Taylor [14]. A list assignment for a graph $G$ is a function $L\colon V(G)\to\operatorname{Pow}(Y)$, where $Y$ is a set, whose elements are referred to as colors. For each $u\in V(G)$, the set $L(u)$ is called the list of $u$ and its elements are said to be available for $u$. A proper coloring $f\colon V(G)\to Y$ is called an $L$-coloring if $f(u)\in L(u)$ for each $u\in V(G)$. A graph $G$ with a list assignment $L$ is said to be $L$-colorable if it admits an $L$-coloring. The list chromatic number $\chi_{\ell}(G)$ of $G$ is the least positive integer $k$ such that $G$ is $L$-colorable whenever $L$ is a list assignment for $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. If $L(u)=\{1,\ldots,k\}$ for a positive integer $k$ for all $u\in V(G)$, then $G$ is $L$-colorable if and only if it is $k$-colorable; in this sense, list coloring generalizes proper coloring. In particular, $\chi_{\ell}(G)\geq\chi(G)$ for all graphs $G$.
The definition of critical graphs can be naturally extended to the list coloring setting. A graph $G$ is list $k$-critical if $\chi_{\ell}(G)\geq k$, but $\chi_{\ell}(G^{\prime})\leq k-1$ for each proper subgraph $G^{\prime}$ of $G$. So, we can define $f_{\ell}(n,k)$ to be the minimum number of edges in an $n$-vertex list $k$-critical graph.
A list assignment $L$ for a graph $G$ is called a degree list assignment if $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$. A fundamental result of Borodin [3] and Erdős, Rubin, and Taylor [14] provides a complete characterization of all graphs that are not $L$-colorable with respect to some degree list assignment $L$. This result can be viewed as a generalization of Theorem 1.2.
@@ 1.2 Known results on list coloring
Let $G$ be a connected graph and $L$ be a degree list assignment for $G$. If $G$ is not $L$-colorable, then $G$ is a Gallai tree; furthermore, $|L(u)|=d_{G}(u)$ for all $u\in V(G)$ and if $u$, $v\in V(G)$ are two adjacent non-cut vertices, then $L(u)=L(v)$.
Exactly as Theorem 1.2 implies (2), Theorem 1.5 yields the same lower bound on $f_{\ell}(n,k)$. Kostochka and Stiebitz [20] improved (2) for $k\geq 9$: the last term in their bound is asymptotically (in $k$) $1.5$ times larger than the one in (2). Then, in a series of papers by Kierstead and Rabern [17], Cranston and Rabern [5], and Rabern [23, 24], this bound was significantly improved and also extended to $4\leq k\leq 8$. The main results by Rabern [23, 24] are as follows:
@@
$f_{\ell}(n,k)\geq\begin{cases}\displaystyle\left(k-1+\frac{(k-3)^{2}(2k-3)}{k^{4}-2k^{3}-11k^{2}+28k-14}\right)\frac{n}{2}&\mbox{ if }k\geq 7\\[8.61108pt] \displaystyle\left(5+\frac{93}{766}\right)\frac{n}{2}&\mbox{ if }k=6\\[8.61108pt] \displaystyle\left(k-1+\frac{k-3}{k^{2}-2k+2}\right)\frac{n}{2}&\mbox{ if }k\in\{4,5\}\\ \end{cases}$
In fact, the bounds were proved not only for $f_{\ell}(n,k)$ but also for the analogous parameter of online list coloring. The last term in Theorem 1.6 is asymptotically (in $k$) twice larger than the one in (2). Below is a table similar to ones in [5, 17, 23] with an added column of our results in this paper.
@@
Table 1: Current lower bounds on the average degree of $k$-critical and list $k$-critical graphs. Best results for list $k$-critical graphs are highlighted.
1.3 Known results on DP-coloring and our results
In this paper we focus on a generalization of list coloring that was recently introduced by Dvořák and Postle [13]; they called it correspondence coloring, and we call it DP-coloring for short. Dvořák and Postle invented DP-coloring in order to approach an open problem about list colorings of planar graphs with no cycles of certain lengths. This hints the usefulness of DP-coloring for graph coloring problems, in particular list coloring problems.
For a multigraph $G$, a (DP-)cover of $G$ is a pair $(H,L)$, where $H$ is a graph and $L\colon V(G)\to\operatorname{Pow}(V(H))$ is a function such that
@@ 1.3 Known results on DP-coloring and our results
We often refer to the vertices of $H$ as colors. A multigraph $G$ with a cover $(H,L)$ has an $(H,L)$-coloring if $H$ has an independent set containing exactly one vertex from $L(v)$ for each $v\in V(G)$. The DP chromatic number $\chi_{\mathrm{DP}}(G)$ of a multigraph $G$ is the least positive integer $k$ such that $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a cover of $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. Every list coloring problem can be represented as a DP-coloring problem. In particular, $\chi_{\mathrm{DP}}(G)\geq\chi_{\ell}(G)$ for all multigraphs $G$.
We say a multigraph $G$ is DP degree-colorable if $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a cover of $G$ with $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$.
A multigraph $G$ is DP $k$-critical if $\chi_{\mathrm{DP}}(G)=k$ and $\chi_{\mathrm{DP}}(G^{\prime})\leq k-1$ for every proper subgraph $G^{\prime}$ of $G$. Let $f_{\mathrm{DP}}(n,k)$ denote the minimum number of edges in an $n$-vertex DP $k$-critical (simple) graph.
For a simple graph $G$ and a positive integer $s$, the multiple $G^{s}$ of $G$ is the multigraph obtained from $G$ by replacing each edge $uv\in E(G)$ with $s$ edges joining $u$ and $v$. In particular, $G^{1}=G$. A GDP-forest is a multigraph in which every block is isomorphic to either $K_{n}^{t}$ or $C_{n}^{t}$ for some $n$ and $t$. (A double cycle is a graph $C^{2}_{n}$.) A GDP-tree is a connected GDP-forest. Note that every Gallai tree is also a GDP-tree.
Theorems 1.2 and 1.5 extend to DP-coloring as follows.
Theorem 1.7 (Dvořák and Postle [13] for simple graphs; Bernshteyn, Kostochka, and Pron’ [2] for multigraphs).
Suppose that $G$ is a connected multigraph. Then $G$ is not DP degree-colorable if and only if $G$ is a GDP-tree.
Analogous to how Theorems 1.2 and 1.5 imply lower bounds on $f(n,k)$ and $f_{\ell}(n,k)$, respectively, Theorem 1.7 yields that (2) holds also for $f_{\mathrm{DP}}(n,k)$ when $n\geq k+2$. Note that the bound (2) does not hold for DP-coloring of general multigraphs: for example, multiple of cycles $C^{k}_{n}$ are $2k$-regular and DP $(2k+1)$-critical.
Bernshteyn and Kostochka [1] proved that for $n\geq 2k$, Dirac’s bound (1) holds also for $f_{\mathrm{DP}}(n,k)$.
@@
Let $k\geq 5$ and $\lambda=\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil$. If $G$ is an $n$-vertex DP $k$-critical graph, then one of the following holds:
  1. 1.
    $G=K_{k}$,
  2. 2.
    $k=5$ and either $G=K^{4}_{2}$ or $G$ is a double cycle, or
  3. 3.
    $|E(G)|\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$
Note that in Case 3 of Theorem 1.8, the average degree of $G$ is greater than
$k-1+\frac{1}{\lambda}=k-1+\left\lceil\frac{k^{2}-7}{2k-7}\right\rceil^{-1}.$
@@
$f_{\ell}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}\mbox{ $\quad$ and $\quad$}f_{\mathrm{DP}}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$ (6)
This is slightly better than the bound of Theorem 1.6 for $k\geq 5$, see the last column of Table 1. We do not know how to prove (6) without using DP-coloring. Thus, this is another example of usefulness of DP-coloring for proving results on list coloring. As it is, the proof of Theorem 1.8 does not work well for $k=4$. In [4], we use a somewhat more elaborate technique to prove that for $n\geq 11$, $f_{\mathrm{DP}}(n,4)>\frac{8}{5}n$.
1.4 Outline of the proof and structure of the paper

### v2

@@ Abstract.
Peter Bradshaw received support from NSF RTG grant DMS-1937241 and an AMS Simons Travel Grant.
Ilkyoo Choi is supported by the Hankuk University of Foreign Studies Research Fund, the Institute for Basic Science (IBS-R029-C1), and the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-23324220).
Alexandr Kostochka is supported in part by NSF Grant DMS-2153507 and by NSF RTG Grant DMS-1937241.
Jingwei Xu is supported in part by Campus Research Board Award RB24000 of the University of Illinois Urbana-Champaign.
1. Introduction
In this paper, a graph has neither parallel edges nor loops. A multigraph may have parallel edges, but no loops. Given a multigraph $G$, let $V(G)$ and $E(G)$ denote the vertex set and edge set, respectively, of $G$. Given a vertex $v\in V(G)$, the degree of $v$, denoted $d_{G}(v)$, is the number of edges incident with $v$. The neighborhood $N_{G}(v)$ of a vertex $v$ is the set of vertices adjacent to $v$. Note that $|N_{G}(v)|\leq d_{G}(v)$, and equality holds if and only if $v$ has no incident parallel edges. A vertex of degree $d$ (at least $d$) is a $d$-vertex ($d^{+}$-vertex). For vertex subsets $S_{1}$ and $S_{2}$, let $E_{G}(S_{1},S_{2})$ denote the set of edges $xy\in E(G)$ where $x\in S_{1}$ and $y\in S_{2}$. For integers $a,b$ with $a\leq b$, let $[a,b]$ denote the set $\{a,a+1,\ldots,b\}$ of integers.
1.1. Known results on proper coloring
A (proper) $k$-coloring of a graph $G$ is a mapping $\varphi\,:\,V(G)\to[1,k]$ such that $\varphi(u)\neq\varphi(v)$ for each $uv\in E(G)$. The chromatic number of $G$, denoted $\chi(G)$, is the minimum positive integer $k$ for which $G$ has a proper $k$-coloring. A graph $G$ is $k$-colorable if $\chi(G)\leq k$. For a positive integer $k$, a graph $G$ is $k$-critical if $\chi(G)=k$, but every proper subgraph of $G$ is (k1)(k-1)-colorable.
The notion of $k$-critical graphs was introduced and systematically studied by Dirac [8, 9, 10, 12, 11, 13] starting from 1951. In particular, Dirac considered the minimum number $f(n,k)$ of edges in an $n$-vertex $k$-critical graph. With this notation, $f(k,k)=\binom{k}{2}$ and $f(k+1,k)$ is not well defined because there is no $(k+1)$-vertex $k$-critical graph. Since every vertex of a $k$-critical graph has degree at least k1k-1, we trivially have $f(n,k)\geq(k-1)\frac{n}{2}$. Even this simple bound yields the Heawood Formula [19] that every graph $G$ embeddable in an orientable surface $S_{\gamma}$ with $\gamma\geq 1$ handles satisfies $\chi(G)\leq\lfloor c_{\gamma}\rfloor$, where $c_{\gamma}=\frac{7+\sqrt{1+48\gamma}}{2}$. The well-known Brooks’ Theorem [5] implies that for $k\geq 4$ and $n\geq k+2$, $f(n,k)\geq\frac{(k-1)n+1}{2}$. Dirac sharpened Heawood’s result by showing that for $\gamma\geq 1$, every graph embeddable in $S_{\gamma}$ with chromatic number $\lfloor c_{\gamma}\rfloor$ contains the complete graph with $\lfloor c_{\gamma}\rfloor$ vertices as a subgraph. For this, he used the following lower bound on $f(n,k)$:
@@ 1.1. Known results on proper coloring
(1) $f(n,k)\geq\left(k-1\right)\frac{n}{2}+\frac{k-3}{2}.$
@@
In his fundamental papers [16] and [17] from 1963, Gallai proved a series of important properties of color-critical graphs. In particular, he found the exact values of $f(n,k)$ for all $n\in[k+2,2k-1]$.
Recall that a block in a graph $G$ is a maximal connected subgraph of $G$ that has no cut vertices. A Gallai tree is a connected graph whose every block is a complete graph or an odd cycle, and a Gallai forest is a graph in which every connected component is a Gallai tree.
@@
Let $k\geq 4$, and let $G$ be a $k$-critical graph. If $B$ is the set of (k1)(k-1)-vertices in $G$, then $G[B]$ is a Gallai forest.
Gallai [16] showed that his Theorem 1.2 implies the following lower bound on $f(n,k)$: if $k\geq 4$ and $n\geq k+2$, then
(2) $f(n,k)\geq\left({k-1}+\frac{k-3}{k^{2}-3}\right)\frac{n}{2}.$
For large $n$, this bound is much stronger than (1).
Krivelevich [27, 28] improved the lower bound on $f(n,k)$ in (2) to
(3) $f(n,k)\geq\left(k-1+\frac{k-3}{k^{2}-2k-1}\right)\frac{n}{2}$
and demonstrated nice applications of his bound. Then, Kostochka and Stiebitz [24] proved that for $k\geq 6$ and $n\geq k+2$,
(4) $f(n,k)\geq\left(k-1+\frac{2(k-3)}{k^{2}+6k-9-6/(k-2)}\right)\frac{n}{2}.$
Note that the last term in the bound (4) is asymptotically (in $k$) twice larger than the one in (2). Then, Kostochka and Yancey [25] proved an asymptotically exact bound: if $k\geq 4$ and $n\geq k,\,n\neq k+1$, then
(5) $f(n,k)\geq\left(k-1+\frac{k-3}{k-1}\right)\frac{n}{2}-\frac{k(k-3)}{2(k-1)}.$
This bound on $f(n,4)$ yields a short proof of the well-known theorem by Grötzsch [18] that every triangle-free planar graph is $3$-colorable.
1.2. Known results on list coloring
List coloring was introduced independently by Vizing [33] and Erdős, Rubin, and Taylor [15]. For a set $X$, let $\operatorname{Pow}(X)$ denote the power set of a set $X$, and denote $\bigcup_{v\in X}f(v)$ by $f(X)$. A list assignment for a graph $G$ is a function $L\colon V(G)\to\operatorname{Pow}(Y)$, where $Y$ is a set, whose elements are referred to as colors. For each $u\in V(G)$, the set $L(u)$ is called the list of $u$ and its elements are said to be available for $u$. A proper coloring $\varphi\colon V(G)\to Y$ is called an $L$-coloring if $\varphi(u)\in L(u)$ for each $u\in V(G)$. A graph $G$ with a list assignment $L$ is said to be $L$-colorable if it admits an $L$-coloring. The list chromatic number $\chi_{\ell}(G)$ of $G$ is the least positive integer $k$ such that $G$ is $L$-colorable whenever $L$ is a list assignment for $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. If $L(u)=[1,k]$ for a positive integer $k$ for all $u\in V(G)$, then $G$ is $L$-colorable if and only if it is $k$-colorable; in this sense, list coloring generalizes proper coloring. In particular, $\chi_{\ell}(G)\geq\chi(G)$ for all graphs $G$.
The definition of critical graphs can be naturally extended to the list coloring setting. A graph $G$ is list $k$-critical if $\chi_{\ell}(G)=k$, but $\chi_{\ell}(G^{\prime})\leq k-1$ for each proper subgraph $G^{\prime}$ of $G$. So, we can define $f_{\ell}(n,k)$ to be the minimum number of edges in an $n$-vertex list $k$-critical graph.
A list assignment $L$ for a graph $G$ is called a degree-list assignment if $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$. A fundamental result of Borodin [3] and Erdős, Rubin, and Taylor [15] provides a complete characterization of all graphs that are not $L$-colorable with respect to some degree-list assignment $L$. This result can be viewed as a generalization of Theorem 1.2.
@@ 1.2. Known results on list coloring
Let $G$ be a connected graph and $L$ be a degree-list assignment for $G$. If $G$ is not $L$-colorable, then $G$ is a Gallai tree; furthermore, $|L(u)|=d_{G}(u)$ for all $u\in V(G)$ and if $u,v\in V(G)$ are two adjacent non-cut vertices, then $L(u)=L(v)$.
Kostochka, Stiebitz, and Wirth [22] and independently Thomassen [32] described for each Gallai tree $G$ the degree-list assignments $L$ such that $G$ is not $L$-colorable. Kostochka and Stiebitz [23] proved that for $n\geq k+2$, Dirac’s bound (1) holds for $f_{\ell}(n,k)$.
Exactly as Theorem 1.2 implies (2), Theorem 1.3 yields the same lower bound on $f_{\ell}(n,k)$. Kostochka and Stiebitz [24] improved (2) for $k\geq 9$: the last term in their bound is asymptotically (in $k$) $1.5$ times larger than the one in (2). Then, in a series of papers by Kierstead and Rabern [20], Cranston and Rabern [6], and Rabern [29, 30], this bound was significantly improved and also extended to $4\leq k\leq 8$. The best known bounds are due to Rabern [29, 30]:
@@
[…]
In fact, the bounds were proved not only for $f_{\ell}(n,k)$ but also for the analogous parameter of paintability, which is defined in terms of an on-line list coloring process where available colors are revealed one at a time (see [29, 30] for a complete definition). The last term in Theorem 1.4 is asymptotically (in $k$) twice larger than the one in (2). Below is a table similar to ones in [6, 20, 29] with an added column of our results in this paper.
@@
Table 1. Current lower bounds on the average degree of $k$-critical and list $k$-critical graphs. Best results for list $k$-critical graphs are marked with an asterisk.
1.3. Known results on DP-coloring and our results
In this paper we focus on a generalization of list coloring that was recently introduced by Dvořák and Postle [14]; they called it correspondence coloring, and we call it DP-coloring for short. Dvořák and Postle invented DP-coloring in order to approach an open problem about list colorings of planar graphs with no cycles of certain lengths. This hints at the usefulness of DP-coloring for graph coloring problems, in particular list coloring problems.
For a multigraph $G$, a (DP-)cover of $G$ is a pair $(H,L)$, where $H$ is a multigraph and $L\colon V(G)\to\operatorname{Pow}(V(H))$ is a function such that
@@ 1.3. Known results on DP-coloring and our results
An $(H,L)$-coloring of $G$ is a function $\phi:V(G)\rightarrow\bigcup_{v\in V(G)}L(v)$ with $\phi(v)\in L(v)$ such that $\{(v,\phi(v)):v\in V(G)\}$ is independent in $H$. In view of this definition, we often refer to the vertices of $H$ as colors. The DP chromatic number $\chi_{\mathrm{DP}}(G)$ of a multigraph $G$ is the least positive integer $k$ such that $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a cover of $G$ with $|L(u)|\geq k$ for all $u\in V(G)$. Every list coloring problem can be represented as a DP-coloring problem. In particular, $\chi_{\mathrm{DP}}(G)\geq\chi_{\ell}(G)$ for all multigraphs $G$.
A cover $(H,L)$ of a multigraph $G$ is called a degree-cover if $|L(u)|\geq d_{G}(u)$ for all $u\in V(G)$. We say a multigraph $G$ is DP degree-colorable if $G$ has an $(H,L)$-coloring whenever $(H,L)$ is a degree-cover.
A multigraph $G$ is DP $k$-critical if $\chi_{\mathrm{DP}}(G)=k$ and $\chi_{\mathrm{DP}}(G^{\prime})\leq k-1$ for every proper subgraph $G^{\prime}$ of $G$. Let $f_{\mathrm{DP}}(n,k)$ denote the minimum number of edges in an $n$-vertex DP $k$-critical graph.
For a graph $G$ and a positive integer $s$, the multiple $sG$ of $G$ is the multigraph obtained from $G$ by replacing each edge $e\in E(G)$ with $s$ edges joining the endpoints of $e$. In particular, $1G=G$. A GDP-forest is a multigraph such that for every block $B$, there exist $n$ and $t$ such that $B$ is either a $tK_{n}$ or a $tC_{n}$. (A double cycle is a multigraph $2C_{n}$.) A GDP-tree is a connected GDP-forest. Note that every Gallai tree is also a GDP-tree. An example of a GDP-tree is shown in Figure 1.
Figure 1. A GDP-tree with two $K_{3}$ blocks, a $2K_{3}$ block, two $K_{2}$ blocks, a $K_{4}$ block, and a $C_{8}$ block.
Theorems 1.2 and 1.3 extend to DP-coloring as follows.
Theorem 1.5 (Dvořák and Postle [14] for graphs; Bernshteyn, Kostochka, and Pron’ [1] for multigraphs).
Let $G$ be a connected multigraph, and let $(H,L)$ be a degree-cover of $G$. If $G$ has no $(H,L)$-coloring, then $G$ is a GDP-tree.
Kim and Ozeki [21] characterized for each GDP-tree the degree-covers $(H,L)$ for which $G$ is not $(H,L)$-colorable.
Analogous to how Theorems 1.2 and 1.3 imply lower bounds on $f(n,k)$ and $f_{\ell}(n,k)$, respectively, Theorem 1.5 yields that the lower bound in (2) holds also for $f_{\mathrm{DP}}(n,k)$ when $n\geq k+2$. Note that the bound (2) does not hold for DP-coloring of general multigraphs: for example, multiples of cycles $kC_{n}$ are $2k$-regular and DP $(2k+1)$-critical.
Bernshteyn and Kostochka [2] proved that for $n\geq k+2$, Dirac’s bound (1) holds also for $f_{\mathrm{DP}}(n,k)$. The reader can find more history of the study of color-critical graphs, including seminal results by Gallai, Hajós, Sachs, and Stiebitz in the recent monographs [7, 31].
@@
(6) $f_{\mathrm{DP}}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$
@@
(7) $f_{\ell}(n,k)\geq\left(k-1+\frac{1}{\lambda}\right)\frac{n}{2}+\frac{1}{\lambda}.$
This is slightly better than the bound of Theorem 1.4 for $k\geq 5$, see the last column of Table 1. We do not know how to prove (7) without using DP-coloring. Thus, this is another example of usefulness of DP-coloring for proving results on list coloring. As it is, an analog of Theorem 1.6 for $k=4$ requires additional ideas to work. In [4], we use a somewhat more elaborate technique to prove that for $n\geq 11$, $f_{\mathrm{DP}}(n,4)>\frac{8}{5}n$, which is better than (6) for $k=4$.
1.4. Outline of the proof and structure of the paper